OpenStudy (anonymous):

Please help me Solve |4x-11| < 5

OpenStudy (saifoo.khan):

Do u have a Million questions?

OpenStudy (saifoo.khan):

haha, then?

OpenStudy (saifoo.khan):

x < 4 x > 3/2

OpenStudy (anonymous):

how did you solve it?

OpenStudy (saifoo.khan):

|4x-11| < 5 4x - 11 < 5 And 4x - 11 > -5

OpenStudy (agreene):

\[3/2<R(x)<4\] and \[ -\frac{\sqrt{-2 R(x)^2+11 R(x)-12)}}{\sqrt{2}}<I(x)<\frac{\sqrt{-2 R(x)^2+11 R(x)-12}}{\sqrt{2}}\] Where R(x) are the reals and I(x) are imaginary parts of Z